OpenStudy (anonymous):

Given that log base 5 of 2 = 0,431 log base 5 of 3 = 0,683 and log base 5 of 7 = 1,209, correct to 3 decimal places find the value of log base 5 of 25

OpenStudy (mathstudent55):

\( \log ab = \log a + \log b\)

OpenStudy (anonymous):

by the exponential interaction of logs \[\log_{5}25=\log_{5}5^{2} =2\log_{5}5=2*1=2 \]

OpenStudy (mathstudent55):

\( \log a^n = n \log a\)

OpenStudy (anonymous):

answer should be to 3 decimal places

OpenStudy (mathstudent55):

2.000

OpenStudy (mathstudent55):

This question doesn't make sense because I don't see how you use log of 2, 3, and 7 to arrive at log of 25.